# Harmonia: A Multi-Agent Reinforcement Learning Approach to Data Placement and Migration in Hybrid Storage Systems

### v1

@@ 6.2. Multi-Programmed Workloads
Figs. 14(a) and 14(b) show the performance of Harmonia and baselines on multi-programmed workloads (see Table 6) on performance- and cost-optimized HSS, respectively. Performance is shown as average I/O latency normalized to Fast-Only.
@@
We make two key observations. First, Harmonia outperforms prior approaches in all multi-programmed workloads across both HSS configurations. In performance-optimized (cost-optimized) HSS, Harmonia (i) improves performance by 38.9% (45.4%), 34.7% (44.8%), 28.4% (33.1%), 32.8% (25.3%), 46.2% (53.0%), 40.3% (47.5%), 30.6% (33.7%), and 39.0% (57.7%) over CDE, K-SVM, RNN-HSS, Sibyl, Sibyl+K-SVM, Sibyl+RNN-HSS, CDE+RL-Migr and SAPM in performance-optimized (cost-optimized) HSS, and (ii) achieves 75% (46%) of Oracle’s performance. Second, Harmonia’s performance improves as the number of concurrent workloads increases. On eight concurrent workloads (e.g., mix6), Harmonia shows slightly higher performance gains, 39.4% (41.8%), over Sibyl compared to other multi-programmed workloads in performance-optimized (cost-optimized) HSS. Harmonia’s RL-based policies learn access pattern variations faster than prior approaches.
We conclude that Harmonia outperforms prior approaches on all multi-programmed workloads due to better adaptability to frequent variations in workload access patterns.
@@ 6.3. Extensibility to Multiple Devices in HSS
We evaluate Harmonia’s extensibility to different HSS configurations by comparing its performance with the best-performing prior approach, Sibyl, on an HSS with three and four devices (see Table 4). We do not evaluate other baselines because they are designed for fixed HSS configuration and lack easy extensibility. Figs. 15(a) and 15(b) show the performance of Sibyl and Harmonia on a Tri-HSS and Quad-HSS respectively. Performance is shown as average request latency normalized to Fast-Only. Harmonia can be easily extended to more devices by adding (1) the remaining capacity of the new storage device to the input features, and (2) a new node in the neural network output layer to represent the new device as a target device for data placement and data migration.
Figure 15. Performance of Sibyl and Harmonia on Tri-HSS (top) and Quad-HSS (bottom) configurations. Performance is shown as average request latency normalized to Fast-Only. Lower is better.
We make two key observations. First, Harmonia outperforms Sibyl by 37.0% (42.0%) on average in Tri-HSS (Quad-HSS). Second, in several workloads (e.g., RocksDB, MLPerf), Harmonia shows higher performance gains (up to 58.3%) over Sibyl than the average performance improvement because Harmonia places performance-critical data in the high-end and mid-range devices and uses its data-migration policy to move cold data to slow devices during idle times. In contrast, Sibyl frequently evicts data to the low-end devices in the critical path, adding significant latency overhead.
We conclude that Harmonia demonstrates extensibility to multiple HSS configurations and achieves high performance across all tested workloads and HSS configurations.
@@ 6.4. Lifetime of Devices in HSS
We evaluate the impact of Harmonia and baselines on storage device lifetimes by measuring their write amplification (WA). WA is the ratio of total data written (including migrations) in the HSS to the total data originating from the workload. A higher WA value adversely impacts the lifetime of storage devices. Figs. 16(a) and 16(b) show WA of Harmonia and baselines on performance-optimized and cost-optimized HSS, respectively.
@@
We make two key observations. First, in both HSS configurations, only Sibyl and RNN-HSS have a lower WA than Harmonia. Harmonia’s average WA of 1.54 (1.55) is higher than Sibyl’s 1.49 (1.45) and RNN-HSS’s 1.08 (1.06) in performance-optimized (cost-optimized) HSS because it proactively migrates data to the fast device in read-intensive workloads to improve read performance. Second, in both HSS configurations, Harmonia has lower WA (1.24) in write-intensive workloads (i.e., SYSTOR17) compared to its WA (1.8) in read-intensive workloads (i.e., RocksDB, MLPerf). In write-intensive workloads, Harmonia learns to move data during update operations and performs fewer migrations.
Write Traffic Distribution in Harmonia. Fig.  17 shows the distribution of writes caused by Harmonia’s migration policy across fast and slow devices in performance-optimized and cost-optimized HSS.
@@
We make two observations. First, Harmonia migrates more data from fast device to slow device in both configurations because it effectively utilizes fast device for frequently-accessed data. Second, Harmonia has more writes to fast device in read-intensive workloads (i.e., RocksDB and MLPerf) than in write-intensive workloads because it proactively migrates frequently-read data to the fast device even when there are no application updates.
We conclude that Harmonia’s adaptive policies provide higher performance at slightly higher average WA compared to the best-performing prior approach.
@@ 7. Discussion
Alternate Objectives and Reward Structures for Harmonia. Reward formulation is critical for finding optimal policies in RL. Careful design of reward structure is necessary to achieve coordination in a multi-agent system with autonomous agents. Harmonia’s focus is to improve HSS performance. Hence, we design the RL agents’ reward structure based on I/O latencies.
However, we can design Harmonia for other objectives such as improving device lifetimes, fairness, quality of service (QoS) and energy efficiency. For example, to improve storage device lifetimes, we can design the state features and reward structure to include (1) number of writes performed on each device, (2) device endurance in terms of program/erase cycles, (3) write amplification factor (WAF), and (4) device reliability properties (e.g., NAND flash memory type (SLC/TLC/QLC), ECC correction capability). To promote or disincentivize write traffic to each device, the reward should be appropriate to the device where the data is placed or migrated to. Harmonia can accommodate fairness and QoS objectives by including factors such as bandwidth utilization, device queue depths and application requirements. Combining different objectives is an interesting research direction, and we leave the exploration of these objectives and reward structures to future work.
Extensibility of Harmonia. We describe Harmonia’s extensibility to a large number of storage devices in the HSS. Harmonia, unlike prior heuristic and supervised learning based approaches, is easily scalable with minimal designer effort. When a new device is added to the HSS, we only need to extend the state features and actions of Harmonia’s RL agents. We add (1) capacity utilization of the new device as a state feature (see Table 2), and (2) an output node in each RL agent’s neural network to represent the action of placing or migrating a page to the new device. However, the larger state and action spaces resulting from scaling Harmonia to more devices leads to two challenges: (1) the data-placement and data-migration policies may take longer to converge to an optimal policy, and (2) the metadata overhead to store the state features increases.
Applicability to Other Systems. Data placement and data migration are critical for the high performance of other heterogeneous systems such as tiered storage systems, hybrid memory systems, disaggregated storage, and networked storage systems. Harmonia can be adapted to these systems, but they come with diverse constraints and requirements.
In a tiered storage system, data placement is always performed on the top-tier device, and hence Harmonia’s data-placement agent can be replaced by a heuristic approach. However, Harmonia’s data-migration agent gains significance due to data migration across tiers. We can add multiple data-migration agents to perform migration between each pair of adjacent tiers.
Hybrid memory systems have strict latency constraints, and require Harmonia to be redesigned for low-latency decision-making. Disaggregated storage systems that use CXL (maruf2023tpp, ; gouk2023memory, ; li2023pond, ) for large-scale memory pooling can leverage Harmonia’s benefits when deployed on CXL fabric controller. Networked storage systems with local and remote storage devices can also benefit from Harmonia’s RL framework. We hope Harmonia can assist in designing adaptive data-management policies for other heterogeneous systems.
@@ 8. Related Work
To our knowledge, Harmonia is the first work to use multi-agent RL for combined optimization of data-placement and data-migration policies in HSS. In this section, we review other related work in HSS data management, and RL in storage/memory systems.
@@
HSS Data Management. A large body of prior work proposes heuristic- and machine-learning-based data-placement (e.g., (matsui2017design, ; sun2013high, ; heuristics_hyrbid_hystor_sc_2011, ; lv2013probabilistic, ; guerra2011cost, ; elnably2012efficient, ; heuristics_usenix_2014, ; bu2012optimization, ; krish2016efficient, ; tai2015sla, ; zhang2010automated, ; wu2012data, ; iliadis2015exaplan, ; lv2013hotness, ; matsui2017tri, ; feng2014hdstore, ; yang2017autotiering, ; singh2022sibyl, ; doudali2019kleio, ; ren2019archivist, ; cheng2019optimizing, ; shetti2019machine, ; li2017utility_HMM, ; agarwal2015page_HMM, ; agarwal2017thermostat_HMM, ; ham2013disintegrated_HMM, ; lin2011hot, ; salkhordeh2015operating, ; pavlovic2013data, ; meza2013case, ; chiachenchou2015batman, ; kim2011hybridstore, ; wang2019panthera, ; ramos2011page, ; liu2019hierarchical, ; luo2020optimal, ; doudali2021cori, ; sen2019machine, )) and data-migration (e.g., (lin2011hot, ; lin2017efficient, ; lin2018buffer, ; zhang2010adaptive, ; cheng2015amc, ; shetti2019machine, ; vengerov2008reinforcement, ; vasilakis2020hybrid2, )) techniques. Hot Random Off-loading (HRO) (lin2011hot, ) uses a 0-1 knapsack model to allocate or migrate files between the storage devices. ADLAM (zhang2010adaptive, ) uses limited workload characteristics to tune data-migration policies in cloud storage systems. Vengerov (vengerov2008reinforcement, ) uses a fuzzy rule base architecture to perform data migration in a hierarchical storage system. These works do not perform (1) real system evaluation or (2) combined optimization of placement and migration policies using multiple RL agents.
@@
Reinforcement Learning in Storage Systems. Several works (e.g., (liu2019learning, ; yoo2020reinforcement, ; wang2020reinforcement, ; rl_GC_TECS, ; kang2018dynamic, ; li2023rlalloc, ; li2020mitigating, ; akgun2021machine, ; kang2017reinforcement, ; kang2019q, ; wei2023reinforcement, ; li2024page, ; liaw2024reinforcement, ; wu2018maximizing, ; ha2024rl, )) use RL to improve different aspects of storage systems. Prior works  (liu2019learning, ; wang2020reinforcement, ) use RL to improve cloud storage utilization. RLAlloc (li2023rlalloc, ) improves SSD throughput and Quality of Service (QoS). Other works (yoo2020reinforcement, ; rl_GC_TECS, ; kang2018dynamic, ; li2020mitigating, ) use RL to optimize SSD maintenance activities (e.g., garbage collection) and error mitigation. Harmonia is orthogonal to all these works and can be used in conjunction with these techniques on an HSS.
RL for Architectural Optimizations. Prior works use RL for data prefetching (e.g., (pythia, ; peled2015semantic, )), memory scheduling (e.g., (ipek2008self, ; multi_scheduler_HPCA_2012, )), cache replacement strategies (liu2020imitation, ) and network-on-chip arbitration (rl_NOC_AIDArc_2018, ; lin2020deepNOC, ). These works provide evidence that RL holds significant potential for intelligent high-performance computer systems.
@@
Multi-Agent RL Systems. Several works propose using multi-agent reinforcement learning (MARL) for various system optimizations (e.g., (rl_opti_TACO_2017, ; qiu2022reinforcement, ; qiu2022simppo, ; qiu2020firm, ; mao2023multi, ; mlsys2023qiu, ; lu2022rlrp, ; jain2017cooperative, ; jain2017coordinated, ; shen2022multiagent, )). Qiu et al. (qiu2022reinforcement, ; qiu2022simppo, ; qiu2020firm, ; mlsys2023qiu, ) use MARL to improve tail latency of Serverless Function-as-a-Service (FaaS), targeting different systems and objectives than Harmonia. Jain et al. (rl_opti_TACO_2017, ; jain2017cooperative, ) use MARL for dynamic partitioning of last level cache and multi-level caches. Shen et al. (shen2022multiagent, ) propose MARL-assisted cache cleaning for shingled managed recording drives. These techniques have four key differences from Harmonia: (1) their objectives are different from HSS data placement and migration, (2) they use identical RL agents unlike Harmonia’s heterogeneous agents, (3) they use joint actions and shared rewards for all agents while Harmonia’s agents have independent actions, and (4) they use table-based RL unlike Harmonia’s neural network based RL. RLRP (lu2022rlrp, ) proposes a deep RL-based replica placement and migration technique for distributed storage systems. RLRP has five key differences from Harmonia: (1) RLRP does not handle typical I/O requests but focuses on replica placement, (2) RLRP’s placement and migration occur at sparse intervals, making it unsuitable for HSS, (3) RLRP retrains its policies only when storage nodes are added or removed, while Harmonia continuously adapts its policies, (4) RLRP considers only a few state features (e.g., capacity utilization), and (5) RLRP’s design has higher storage overhead than Harmonia. We conclude that these reasons make RLRP unsuitable for data placement and migration in HSS.

### v2

@@ 6.2. Multi-Programmed Workloads
Figs. 13(a) and 13(b) show the performance of Harmonia and baselines on multi-programmed workloads (see Table 5) on performance- and cost-optimized HSS, respectively.
@@
We make two key observations. First, Harmonia outperforms prior approaches in all multi-programmed workloads across both HSS configurations. In performance- (cost-) optimized HSS, Harmonia (i) outperforms Sibyl by 32.8% (25.3%), and (ii) achieves 75% (46%) of Oracle’s performance. Second, Harmonia’s performance improves as the number of concurrent workloads increases. On eight concurrent workloads (e.g., mix6), Harmonia shows higher performance gains, 39.4% (41.8%), over Sibyl compared to other multi-programmed workloads. Harmonia’s RL-based policies learn access pattern variations faster than prior approaches.
@@ 6.3. Extensibility to Multiple Devices in HSS
We evaluate Harmonia’s extensibility to different HSS configurations by comparing its performance with the best-performing prior approach, Sibyl, on an HSS with three and four devices (see Table 3). We do not evaluate other baselines as they are designed for fixed HSS configurations and lack easy extensibility. Figs. 14(a) and 14(b) show the performance of Sibyl and Harmonia on a Tri- and Quad-HSS. We discuss Harmonia’s extensibility in §7.
Figure 14. Performance of Sibyl and Harmonia on Tri-HSS (top) and Quad-HSS (bottom). Performance is shown as average request latency normalized to Fast-Only. Lower is better.
We make two key observations. First, Harmonia outperforms Sibyl by 37.0% (42.0%) on average in Tri-HSS (Quad-HSS) as Harmonia places performance-critical data in the high-end and mid-range devices and uses its migration policy to move cold data to slow devices during idle times. In contrast, Sibyl frequently evicts data to the low-end devices on the critical path, adding significant latency overhead. Second, in read-intensive workloads (e.g., RocksDB, MLPerf), Harmonia shows higher performance gains (up to 58.3%) over Sibyl because Sibyl lacks a migration policy and relies on application updates to migrate data across devices.
@@ 6.4. Lifetime of Devices in HSS
We evaluate the impact of Harmonia and baselines on storage device lifetimes by measuring their write amplification (WA). WA is the ratio of total data written (including migrations) in the HSS to data originating from the workload. A higher WA adversely impacts device lifetimes. Figs. 15(a) and 15(b) show WA of Harmonia and baselines on performance- and cost-optimized HSS, respectively.
@@
We make two key observations. First, only Sibyl and RNN-HSS have a lower WA than Harmonia. Harmonia’s average WA of 1.54 (1.55) is higher than Sibyl’s 1.49 (1.45) and RNN-HSS’s 1.08 (1.06) in performance- (cost-) optimized HSS as it proactively prefetches data to the fast device in read-intensive workloads to improve read performance. Second, in both HSS configurations, Harmonia has lower WA (1.24) in write-intensive workloads (i.e., SYSTOR17) compared to its WA (1.8) in read-intensive workloads (i.e., RocksDB, MLPerf). In write-intensive workloads, Harmonia learns to move data during update operations and performs fewer migrations.
Write Traffic Distribution in Harmonia. Fig.  16 shows the write distribution of Harmonia’s migration policy across devices in performance- and cost-optimized HSS.
@@
We make two observations. First, Harmonia migrates more data to slow device as it effectively utilizes fast device for frequently-accessed data. Second, Harmonia has more writes to fast device in read-intensive workloads (i.e., RocksDB, MLPerf) as it proactively prefetches frequently-read data to the fast device even when there are no application updates.
We conclude that Harmonia’s adaptive policies provide higher performance at slightly higher average WA compared to Sibyl.
@@ 7. Discussion
Alternate Objectives and Reward Structures. Since Harmonia’s focus is to improve HSS performance, we design the RL agents’ reward structure based on I/O latencies. We can design Harmonia for other objectives such as device lifetimes, fairness, quality of service (QoS) and energy efficiency. For example, to improve device lifetimes, the state features and reward structure can include (1) writes performed on each device, (2) device endurance in terms of program/erase cycles, (3) write amplification (WA), and (4) device reliability properties (e.g., NAND flash type (SLC/TLC/QLC), ECC capability). To manage write traffic to each device, the reward should be appropriate to the device where the data is placed or migrated. Harmonia can accommodate fairness and QoS objectives by including factors such as bandwidth utilization, device queue depths and application requirements. We leave the exploration of these objectives and reward structures to future work.
Extensibility of Harmonia. Harmonia is easily extensible to a large number of storage devices with minimal designer effort. When a new device is added to the HSS, we only add (1) capacity utilization of the new device as a state feature (see Table 1), and (2) an output node in each RL agent’s neural network to represent the action of placing or migrating a page to the new device. However, scaling Harmonia to more devices can lead to two challenges: (1) the large state and action spaces can delay the convergence to an optimal policy, and (2) the metadata overhead increases.
Applicability to Other Systems. Harmonia can be adapted to other heterogeneous systems such as tiered storage, hybrid memory, disaggregated storage and networked storage, which have diverse constraints and requirements. In a tiered storage system, placement is always performed on the top-tier device, and hence Harmonia’s placement agent can be replaced by a heuristic approach. The migration agent gains significance due to data migration across tiers. We can add multiple migration agents for each pair of adjacent tiers. Hybrid memory systems have strict latency constraints, and require Harmonia to be redesigned for low-latency decision-making. Disaggregated storage systems that use CXL (maruf2023tpp, ; gouk2023memory, ; li2023pond, ) for large-scale memory pooling can leverage Harmonia’s benefits when deployed on CXL fabric controller. We hope Harmonia can assist in designing adaptive policies for other heterogeneous systems.
@@ 8. Related Work
In this section, we review other related work in HSS data management, and RL in storage/memory systems.
@@
HSS Data Management. A large body of prior work proposes heuristic- and machine-learning-based data-placement (e.g., (matsui2017design, ; sun2013high, ; heuristics_hyrbid_hystor_sc_2011, ; lv2013probabilistic, ; guerra2011cost, ; elnably2012efficient, ; heuristics_usenix_2014, ; bu2012optimization, ; krish2016efficient, ; tai2015sla, ; zhang2010automated, ; wu2012data, ; iliadis2015exaplan, ; lv2013hotness, ; matsui2017tri, ; feng2014hdstore, ; yang2017autotiering, ; singh2022sibyl, ; doudali2019kleio, ; ren2019archivist, ; cheng2019optimizing, ; shetti2019machine, ; li2017utility_HMM, ; agarwal2015page_HMM, ; agarwal2017thermostat_HMM, ; ham2013disintegrated_HMM, ; lin2011hot, ; salkhordeh2015operating, ; pavlovic2013data, ; meza2013case, ; chiachenchou2015batman, ; kim2011hybridstore, ; wang2019panthera, ; ramos2011page, ; liu2019hierarchical, ; luo2020optimal, ; doudali2021cori, ; sen2019machine, )) and data-migration (e.g., (lin2011hot, ; lin2017efficient, ; lin2018buffer, ; zhang2010adaptive, ; cheng2015amc, ; shetti2019machine, ; vengerov2008reinforcement, ; vasilakis2020hybrid2, )) techniques. Hot Random Off-loading (HRO) (lin2011hot, ) uses a 0-1 knapsack model to allocate or migrate files between the storage devices. Vengerov (vengerov2008reinforcement, ) uses a fuzzy rule base architecture to perform data migration in a hierarchical storage system. These works do not perform (1) real system evaluation or (2) combined optimization of placement and migration policies using multiple RL agents.
@@
RL in Storage Systems. Several works (e.g., (liu2019learning, ; yoo2020reinforcement, ; wang2020reinforcement, ; rl_GC_TECS, ; kang2018dynamic, ; li2023rlalloc, ; li2020mitigating, ; akgun2021machine, ; kang2017reinforcement, ; kang2019q, ; wei2023reinforcement, ; li2024page, ; liaw2024reinforcement, ; wu2018maximizing, ; ha2024rl, )) use RL to improve different aspects of storage systems. Prior works  (liu2019learning, ; wang2020reinforcement, ) use RL to improve cloud storage utilization. RLAlloc (li2023rlalloc, ) improves SSD throughput and Quality of Service (QoS). Other works (yoo2020reinforcement, ; rl_GC_TECS, ; kang2018dynamic, ; li2020mitigating, ) use RL to optimize SSD maintenance activities (e.g., garbage collection) and error mitigation. Harmonia is orthogonal to all these works and can be used in conjunction with these techniques on an HSS.
@@
Multi-Agent RL Systems. Several works propose using multi-agent RL (MARL) for various system optimizations (e.g., (rl_opti_TACO_2017, ; qiu2022reinforcement, ; qiu2022simppo, ; qiu2020firm, ; mao2023multi, ; mlsys2023qiu, ; lu2022rlrp, ; jain2017cooperative, ; jain2017coordinated, ; shen2022multiagent, )). Qiu et al. (qiu2022reinforcement, ; qiu2022simppo, ; qiu2020firm, ; mlsys2023qiu, ) use MARL to improve tail latency of Serverless Function-as-a-Service (FaaS), targeting different systems and objectives than Harmonia. Jain et al. (rl_opti_TACO_2017, ; jain2017cooperative, ) use MARL for dynamic partitioning of last level cache and multi-level caches. Shen et al. (shen2022multiagent, ) propose MARL-assisted cache cleaning for shingled managed recording drives. These techniques have four key differences from Harmonia: (1) their objectives are different from HSS data placement and migration, (2) they use identical RL agents unlike Harmonia’s heterogeneous agents, (3) they use joint actions and shared rewards for all agents, and (4) they use table-based RL unlike Harmonia’s neural network based RL. RLRP (lu2022rlrp, ) proposes a deep RL-based replica placement and migration technique for distributed storage systems. RLRP has five key differences from Harmonia: (1) RLRP does not handle typical I/O requests but focuses on replica placement, (2) RLRP’s placement and migration occur at sparse intervals, making it unsuitable for HSS, (3) RLRP retrains its policies only when storage nodes are added or removed, while Harmonia continuously adapts its policies, (4) RLRP considers only a few state features (e.g., capacity utilization), and (5) RLRP’s design has higher storage overhead than Harmonia. We conclude that these reasons make RLRP unsuitable for data placement and migration in HSS.